satellite73 (satellite73):

Adam and Eve roll a die. if it shows 1 or 2, Eve wins an apple from Adam and if it shows 3, 4, 5, 6 Adam wins an apple from Eve. Eve has two apples, Adam has only one. What is the probability Adam gets both Eve's apples before he loses his.

OpenStudy (bahrom7893):

The SNAKE WINS!

OpenStudy (bahrom7893):

Okay, lets try this (though I suck at probability): P(A1) => adam wins apple 1 P(A2|A1) => adam wins apple 2 given that he won apple 1.

OpenStudy (bahrom7893):

oh crap.. there's also before he loses his. Does finding the probability of Adam winning both apples help?

OpenStudy (bahrom7893):

P(A1) = 4/6=2/3

OpenStudy (anonymous):

yes it does help to figure out what the probability he wins right away. that will get to the solution

OpenStudy (bahrom7893):

P(A2) = 4/6 = 2/3, so P(A2|A1) = 4/9/2/3 = 2/3 ?

OpenStudy (bahrom7893):

I'm still shaky on P(A2|A1) formula. If you want to find probability of A|B do u add or multiply?

OpenStudy (bahrom7893):

for numerator?

OpenStudy (anonymous):

don't worry about conditional probability formula. trials are independent, so think about how adam can win quickest, and then what happens if he doesn't win in two rolls

OpenStudy (bahrom7893):

oh yea i just realized it.

OpenStudy (bahrom7893):

so P(Adam wins quickest) = 4/9 P(Adam loses on first) = P(Adam loses on second) = 1/3

OpenStudy (bahrom7893):

do i just subtract now?

OpenStudy (anonymous):

adam can't lose on first roll and still win the game, so ignore that one.

OpenStudy (bahrom7893):

oh wait so that's it? 4/9?

OpenStudy (anonymous):

either wins on first two tosses, or then wins and loses. what happens then?

OpenStudy (bahrom7893):

4/9 - 1/3?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

not 2/3

OpenStudy (anonymous):

8/9